OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the substitution method. x - y = 0 3x + y = -4

OpenStudy (abbycross167):

My answer~ Solve one equation for either variable~ Y=0+x In the other equation replace the variable from step 1 with what is equaled: 3x+(0+x)=-4 Solve the equation in step 2 for the other variable: 4x= -4 --- x= 4/-4 = -1 Substitute the value found in step 3 into equation in step 1: Y=0+-1 Solve equation in step 4: Y=-1 Write your answer as an ordered pair (x, y). (4/-4, 0/-1)

OpenStudy (abbycross167):

What my teacher said was wrong~when you move the x in the 1st equation, it isn't +x

OpenStudy (alexandrea.lane):

When you move the x in the 1st equation it is -x but you would divide -1 from the y so you would get positive x...

OpenStudy (abbycross167):

So change this.... Solve one equation for either variable~ Y=0+x into this... Solve one equation for either variable~ Y=0-x?

OpenStudy (alexandrea.lane):

no. You should have y=(x+0)

OpenStudy (abbycross167):

oh ok, so change it into Solve one equation for either variable~ Y=(x+0)

OpenStudy (alexandrea.lane):

ya

OpenStudy (abbycross167):

ok then what?

OpenStudy (alexandrea.lane):

do you understand?

OpenStudy (abbycross167):

somewhat... The first 3 steps I'm little confused with

OpenStudy (abbycross167):

what is the -x -x for?

OpenStudy (alexandrea.lane):

Just showing i was subtracting -1x

OpenStudy (abbycross167):

Oh ok, if I write out the steps will you check to make sure I wrote them out right?

OpenStudy (abbycross167):

@alexandrea.lane

OpenStudy (abbycross167):

Solve one equation for either variable~ x- y= 0 -x -x -y=-x+0 -1 -1 Y=x+0 In the other equation replace the variable from step 1 with what is equaled: 3x+x-0=-4 4x -0= -4 4x= -4 X= -1 Substitute the value found in step 3 into equation in step 1: -1-y=0 Solve equation in step 4: Y=-1 Write your answer as an ordered pair (x, y). (-1,-1)

OpenStudy (alexandrea.lane):

It should be -0

OpenStudy (abbycross167):

what part should be?

OpenStudy (alexandrea.lane):

y=x-0